Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

13 buttons in a bag 9 are white 4 are black if you take two buttons out what is the probability that you get pne of each colour

OpenStudy (anonymous):

With or without replacement?

OpenStudy (anonymous):

With replacement, you have a 9/13 chance of getting a white button, and a 4/13 chance of getting a black button. The probability of getting a black and a white is equal two the product of the two probabilities: 9/13*4/13=36/169 or about 21.3%

OpenStudy (anonymous):

Without replacement it gets a bit more complicated... There are now two ways you can get one of each: by choosing a white then a black, or by choosing a black then a white. since there is no replacement, its a changing fraction problem (you are drawing from a different number of buttons on the second draw) Probability of choosing one of each = probability of black then white + probability of white then black = 4/13 * 9/12 + 9/13 * 4/12 = 6/13 or about 46%

OpenStudy (anonymous):

In both cases you need to use \({2 \choose 1}\) to account for black then white, and white then black.

OpenStudy (anonymous):

Notice that \(9\times 4 + 9\times 9+4\times 4 \neq 13\times 13\)

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!